Two charges of value 2micro c and negative 5o micro c are placed at adistance 80 cm apart. calculate the distance point from the smaller charge where the field intensity will be zero
 
 
 
 

Grade:12

1 Answers

Arun
6 years ago
Dear Priya
 
Let the diatance from smaller charge = x
Hence
K* 2 /x² + K* -50/(80+x)² = 0
(80+x)²/x² = 25 = 5²
80+x = 5x
4x = 80
x = 20 cm
 
Regards
Arun (askIITians forum expert)